Solve the following equation for 'x'.
2(x+1)^2+x=-1

Answer by checkley71(8403)

2(X+1)^2+X=-1
2(X^2+2X+1)+X=-1
2X^2+4X+2+X=-1
2X^2+5X+2+1=0
2X^2-5X+3=0
(2X-3)(X-1)=0
2X-3=0
2X=3
X=3/2 ANSWER.
X-1=0
X=1 ANSWER.